# Simultaneous analysis of approximate leave-one-out cross-validation and mean-field inference

### v1

@@ 5. Conditioning example II: Single-index models
In robust linear regression with (4.1)-(4.2), there is only direction $\bar{h}\in\mathbb{R}^{p}$ that we need to condition upon on the right of $X$ in (4.4), because in this event (and for fixed $\varepsilon$), $(y_{i},L_{y_{i}},x_{i}^{T}(\hat{b}-\beta^{*}),y_{i},\hat{b}-\beta^{*},\hat{b})$ are all deterministic. In single index model, the nonlinearity requires us to condition on another direction on the right of $X$.
@@
Let $\mu,\delta>0$ be constants. Assume the proportional regime (1.3). Consider the single index model (5.1) with $\varepsilon$ independent of $X$, and the estimator $\hat{b}$ in (1.1). Assume that $X$ has iid $N(0,\Sigma)$ rows with $\Sigma$ invertible. Assume that for all $y\in\mathcal{Y}$, the function $L_{y}(\cdot)$ is differentiable with $L_{y}^{\prime}(\cdot)$ 1-Lipschitz. Consider one of the regularizer in Examples 1.1, 1.2 and 1.3 and assume $\nu\|\Sigma\|_{op}\geq\mu$ so that the strong-convexity (1.10) holds. For the explicit matrix $\hat{A}$ given in Examples 1.1, 1.2 and 1.3 we have
@@
The proof is exactly the same as in Corollary 4.1 with the following difference: Here we choose choose $P,Q$ in Theorem 3.1 be projections of rank $p-2$ and $n-1$ such that the event (3.1) is the same as (5.2).
@@
where $r>1,q>1$ and $11/r+1/q=$ 1 and $r=2q$, which gives $(1/2+1)/q=1$ so $q=3/2$ and $r=3$. By convexity of $\phi$ and an average of 3, we have
@@
:-$\coloneq\tilde{\mathbb{E}}[A_{i}+B_{i}+C_{i}+E_{i}].:$
@@
Since $X-QXP=(I_{n}-Q)X+QX(I_{p}-P)$, the rank of $X-QXP$ is at most $d_{P}+d_{Q}$ hence

### v2

@@ 5. Conditioning example II: Single-index models
In robust linear regression with (4.1)-(4.2), there is only direction $\bar{h}\in\mathbb{R}^{p}$ that we need to condition upon on the right of $X$ in (4.4), because in this event (and for fixed $\varepsilon$$)(y_{i},L_{y_{i}},x_{i}^{T}(\hat{b}-\beta^{*}),y_{i},\hat{b}-\beta^{*},\hat{b}$) are all deterministic. In single index model, the nonlinearity requires us to condition on another direction on the right of $X$.
@@
Let $\mu,\delta>0$ be constants. Assume the proportional regime (1.3). Consider the single index model (5.1) with $\varepsilon$ independent of $X$, and the estimator $\hat{b}$ in (1.1). Assume that $X$ has iid $N(0,\Sigma)$ rows with $\Sigma$ invertible. Assume that for all $y\in\mathcal{Y}$, the function $L_{y}(\cdot)$ is differentiable with $L_{y}^{\prime}(\cdot)$ 1-Lipschitz. Consider one of the regularizer in Examples 1.1, 1.2 and 1.3 and assume $\nu\|\Sigma\|_{op}\geq\mu$ so that the strong convexity (1.10) holds. For the explicit matrix $\hat{A}$ given in Examples 1.1, 1.2 and 1.3 we have
@@
The proof is exactly the same as in Corollary 4.1 with the following difference: Here we choose choose $P,Q$ in Theorem 3.1 be projections of rank p2p-2 and n1n-1 such that the event (3.1) is the same as (5.2).
@@
where $r>1,q>1$ and $1/r+1/q=1$ and $r=2q$, which gives $(1/2+1)/q=1$ so $q=3/2$ and $r=3$. By convexity of $\phi$ and an average of 3, we have
@@
$\coloneq\tilde{\mathbb{E}}[A_{i}+B_{i}+C_{i}+E_{i}].$
@@
Since $X-QXP=(I_{n}-Q)X+QX(I_{p}-P)$, the rank of XQXPX-QXP is at most $d_{P}+d_{Q}$ hence